OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)?

hero (hero):

10/2 = 5

OpenStudy (anonymous):

Slope = Rise/ Run Slope = y1-y2 / x1-x2 (4,14) - > x1 = 4, y1 = 14 (2,4) -> x2 = 2, y2 = 4 Slope = (14-4) / (4-2) Slope = 10/2 Slope = 5

OpenStudy (anonymous):

thank you both! (:

OpenStudy (anonymous):

Yw,

OpenStudy (karatechopper):

5 is your slope